# A Topological Gaussian Mixture Model for Bone Marrow Morphology in Leukaemia Qiquan Wang1,†, Anna Song1,2, Antoniana Batsivari2, Dominique Bonnet2, and Anthea Monod1

### v1

@@ 4.2 Spatial Texture Decomposition via Local Texture Analysis
In what follows, we will study a texture decomposition (see, e.g., Figure 8) and thus refer to the three groups of textures coloured in white, light blue, and dark blue as textures A, B, and C, arbitrarily, as previously discussed.
We first perform a local texture analysis using the measures and critical sizes as defined in Section 4.1, we seek to analyze the localized textures by computing “local” SDPH diagrams at equidistant locations throughout the 3D image sample for each bone and determining how the distribution of textures on the bone differ with respect to the progression of the disease.
We consider the textural makeup of the knee region and the long region of the bones separately. We select 3D balls spaced at equal distance, each with centers five voxels apart and with radii of 40 voxels in the horizontal plane and 1/5 of the sample thickness in the vertical plane. This contributes to 22 million such balls for the entire dataset of all 27 bones. To facilitate the computation of SDPH diagrams from the bone images, we divide the images into 12 subregions for the long bone and 8 subregions for the knee bone, and compute the SDPH diagrams over these subregions separately. Then we restrict the SDPH diagrams to individual balls by including only points for which both critical points lie within the ball. For each grid point, we consider the critical points captured in its local neighborhood to be representative and to quantify the local morphology at those points. Points with persistence less than 0.5 (approximately less than 0.25m) are removed, since these points of low persistence can be associated with noise in the images when considering their relation to the size of the vessels, which are of the order of 10–20m.
@@
After normalizing the localized feature vectors, we further group these feature vectors into clusters comprising similar textures. Our dominant approach utilizes the -means algorithm (MacQueen et al.,, 1967) in clustering the locations into groups; the -means algorithm is an iterative method to partition the dataset into clusters of similar features. By initializing on some random centroids, then iteratively assigning data points to their nearest centroid, then updating the centroids with the means of the assigned data points, we converge to stable centroids and cluster assignments for the data. We also considered other approaches using alternative clustering techniques, such as Gaussian mixture model clustering (Reynolds,, 2009) and clustering for large applications (CLARA, Kaufman & Rousseeuw, (1990)) which yielded similar results. These results can be found in Appendix B.
@@
Figures 7 and 17 show the distributions of groups of different textures in the knee and long regions of the bone for the different samples when the number of clusters is set to , represented in the figures by the different colors projected back onto their locations in the bone image. From Figure 7, we observe that the CTRL samples consist of largely texture A featured in white, whereas other samples with higher levels of engraftment contain more of the other two textures B and C, shown by the dark and light blue colors. We also see a transition in the ratio of different colored textures across the samples. Figure 8 shows the varying percentage decomposition of the three different textures for the knee region of each bone sample.
@@
Figure 9 shows the percentage decomposition of the three textures projected onto two dimensions for visualization using principal component analysis (PCA), a dimensionality reduction technique which seeks to emphasize the significant features of the data while retaining as much variation of the data as possible. Based on this, we can approximately divide the samples into three groups, which we can see highlighted in the figure in blue, orange, and yellow. It is important to note that this grouping is not unique to this specific combination of and -means clustering algorithm; we find that across a range of values of and both the GMM and CLARA clustering algorithms, we repeatedly observe clear divisions between the outlined Phase I and II samples. On the other hand, Phase O samples may appear distinct to the other two groups or merged with the other groups for some combinations of and algorithms, however the occurrence of the latter is the rarer of the two. Certain samples may violate the grouping, for example, for certain combinations, sample P1 (40%) may appear closer to Phase I samples than Phase II samples. Further evidence is given in Appendix B.
@@
Overall, based on the percentage texture decomposition, we allocate the knee samples into three morphological stages—Phases O, I, and II—as follows:
@@
Returning to Figure 8, we find that, as previously noted, the dominant texture for Phase O samples is texture A, constituting on average over 60% of the texture decomposition—a considerably higher proportion than that found within the other phases. In contrast, for Phase I samples, the dominance shifts from texture A to texture B, with the samples containing on average between 45% to 55% of texture B. Lastly, for Phase II samples, we observe on average a more balanced distribution between the textures, with a similar proportion for each texture. Morphologically, texture A is characterized by thicker vessels compared to textures B and C, whereas texture B features lower values than the other textures and smaller interstitial gaps between the vessels. In contrast, texture C features much larger interstitial gaps between the vessels than found in the other two textures. Further analysis of individual textures can be found in Song, (2023).
Furthermore, we find a correspondence between the phase allocations and the levels of engraftment. By assignment, we imposed that Phase O be solely associated with the CTRL samples (in the “control” stage with no engraftment), while Phase I is associated with samples with low percentages of engraftment, mostly varying between 1% and 25% of engraftment; we describe this as the early stages of engraftment. In comparison, Phase II contains samples with higher percentages of engraftment, varying between 40% and 90%; we associate this with late stages of engraftment. The exceptions to this trend are the two samples MNC at 53% and P1 at 67%, whose actual levels of engraftment may be inaccurately estimated since the values are estimated based on crushing a mixture of bones other than the femur bones being imaged.
Applying the same decomposition on the long region of the samples, however, does not yield similar results. We found that in the long region, parts of the vasculature at late stages of engraftment become similar to the CTRL samples.
@@ 4.3 Spatial Texture via Global Texture Analysis
In parallel, we performed a global analysis of the textures by computing SDPH on the entire knee and long regions of the bone samples and observed commonalities with the phase categories from local texture analysis of the bone samples. To visualize the large collection of birth–death points in persistence diagrams computed for each sample, we compute an approximate density as a Gaussian kernel density estimate. In particular, on each point in the diagram we applied the Gaussian kernel, with , weighted by their persistence which is computed as the difference between the birth and death times. Figure 10 shows an example of the computed Gaussian kernel density estimates plotted as a heatmap with contouring for the SDPH diagrams of the knee region of the first CTRL sample.
@@
We observe several different types of transitions in density in the different quadrants of the persistence diagrams as the samples vary from different types and levels of engraftments. In particular, Table 3 shows the persistence-weighted heatmaps of the density estimates for the PH1 NW quadrants of the SDPH diagrams of the knee samples. We observe increasing variation in values for higher levels of engraftment and from Phase I samples to Phase II samples. This variation suggests that the vascular loops dilate towards later stages of engraftment. Also, moving from Phase O to Phase II samples, we observe a shift in the range of values towards lower values of , suggesting thinner vessels and perhaps thinner vessel walls and greater permeability related to later stages of AML. The same can be observed from samples from the long region of the bone as shown in Table 4, although less prominent.
@@
Another interesting feature is the spread of the points in the PH2 NE quadrant of the persistence diagrams. We note from Tables 2 and 5 that in general, the samples with higher levels of engraftment show greater variation in the distribution of the points in the direction parallel to the diagonal. Other changes can be found in the other quadrants similarly relating the level of engraftment and phase.
Phase O | Phase I | Phase II | |||
Phase O | Phase I | Phase II | |||
To quantify these differences, we take the pairwise norm between each density and construct a rooted tree using the unweighted pair group method with arithmetic mean (UPGMA) (Sokal et al.,, 1958) to predict the global clustering behavior of the samples based on the similarities between the densities. We recall that UPGMA is an agglomerative hierarchical clustering method, which begins with each data sample as a unique cluster and at each step merges the two closest clusters and updates the intergroup distances by taking the mean of the distances between the elements of each group. With the acquired clustering, we detect several strong agreements in the clustering with the local texture analysis from Section 4.2, most prominently within the PH1 and PH2 trees and clustering. We can infer from the tree computed from PH1 persistence-weighted diagrams shown in Figure 26 that the collection of knee region samples can be distinctly separated into two groups of Phase II and Phase O/I samples by taking a slice at 0.1 branch length (except for the anomaly of one P1 sample with 40% engraftment). The branch of Phase O/I samples further divides into two groups, one containing Phase I U937 and HL60 samples mostly with high levels of engraftment relative to its other samples of the same type, and one containing a mix of CTRL samples and samples with low levels of engraftment relative to other samples of the same type.
Analogously, for PH2 in Figure 27, slicing of the dendrogram at 0.22 along the branch length, we obtain three large clusters: one exclusively consisting of Phase II samples; one containing mostly Phase I samples with one CTRL sample; and a smaller cluster consisting of three CTRL samples, the MNC (53%) and P1 (40%) samples, which where deemed more similar to the three CTRL samples by the metric.
Specifically, a region that most clearly reflects the phase classification from the local texture analysis is the PH1 NW quadrant, which indicates the number and size of the loops in the sample images. We see in Figure 11 that the tree first divided into two clusters of Phase II and Phase O/I (again, with the anomaly of the P1 sample with 40% engraftment belonging to the Phase O/I cluster). Then the Phase O/I branch divides into a group of Phase I samples with higher levels of engraftment and a group of Phase O/I samples with lower levels of engraftment, which then in turn is decomposed exactly into a group of CTRL samples (Phase O) and a group of Phase I samples. This alternative global analysis mirrors the results of the local analysis.
@@ 5 Modeling Textural Changes with Gaussian Mixture Models on SDPH Diagrams
In Section 4, we have shown that associations can be made between the SDPH diagrams and morphological features in the bone samples. This indicates that SDPH diagrams can quantify structural changes in bone marrow vasculature in an interpretable manner that is consistent with biological reasoning. Using localized analysis of the images, we can assign the knee region samples into one of three morphologically-derived phases, O, I, and II, which approximately correlates to the control, early and late stages of engraftment. Further, when computing the global SDPH over the entire knee regions of the bone samples and approximating the point distributions using a Gaussian kernel density estimate, we confirmed the concordance with the phase allocations. We identified certain transitions in the density patterns within the SDPH diagrams as we move across the different phases towards progressively higher levels of engraftment. To capture this behavior, we turn to our second contribution of this paper and we now propose a family of phase-dependent Gaussian mixture models (GMMs) to model the common distribution observed for samples from each phase.
For the remainder of this paper, we will focus on studying the SDPH diagrams computed on the knee regions of the samples where the phase allocations are more in line with the progression of the disease. Whenever we refer to the samples, we implicitly imply the knee regions of the samples only. Our reasoning for this restriction is that the nonlinear spread of AML along the femur bone gives rise to mixtures of varying morphology corresponding to contrasting levels of engraftment. We hypothesize that in the early stages, this sporadic behavior is due the non-homogeneity of leukaemic engraftment with the foci of leukaemia present in certain regions but not others, and the tendency for leukaemia engraftment to begin in the epiphysis. Then in the later stages, we speculate this sporadic behavior to persist, as not all blood vessels are being altered by leukaemic cells the same and distribution of damaged vessels will be non-uniform. We observe that at later stages of engraftment, in the long region, the vasculature seems to revert to a more control-like behavior, whereas in the early stages, the proportion of textures is more distinct from the control state. We suspect that rather than the vessels returning control-like structure, this change is in fact the result of space constraints lowering angiogenesis, AML invading the bone marrow, and vascular damage. Hence, whenever the SDPH of the long bone region becomes relevant, it will be explicitly stated.
@@ 5.1 Gaussian Mixture Models
Gaussian mixture models have been widely used for unsupervised clustering and density estimation (Reynolds,, 2009); the models assume the data to be generated from a weighted mixture of Gaussian distributions which are parameterized by some unknown means and variances. Where the -means algorithm is considered a hard clustering technique, GMMs allow for soft clustering, where cluster assignments are associated with probabilities, allowing the model to capture more complex patterns in the data distribution. It is also common for the parameters of the GMM to depend on external parameters, such as time for applications in finance (Zhang et al.,, 2020) and engineering (Mas’ud et al.,, 2021), allowing for context to be integrated into the model.
Due to the discrete nature of our samples and uncertainty in terms of the levels of engraftment, we introduce instead a family of phase-dependent GMMs to approximate our samples’ SDPH diagrams. For each point in the SDPH diagram for each sample, we relabel the indices, so that points from samples of the same phase are viewed as a collective: Let for denote the collection of birth–death points from the SDPH diagrams of samples within phase , where denotes the number of points in this phase and references the individual points in this collection. In addition, each point is weighted by its persistence computed as . Let be the size of the mixture for each phase, then the models are given by
@@
Expectation–Maximization (EM) Algorithm.
To estimate the parameters , we use an iterative expectation–maximization (EM) algorithm, which can be broken down into two steps: An E-step, computing the probability of belonging to each Gaussian distribution,
@@
We determine the size of the mixture model through trade-offs between the model performance evaluated using the Bayesian information criterion (BIC), the computational time, and the generalizability of the model across SDPH diagrams of all samples in the same phase. BIC is chosen for the application since it poses heavier penalties on the model complexity compared to other criteria, such as the Akaike information criterion (AIC) and is ideal for our model.
@@
Since we have no further way of generating more images from each engrafted sample or obtaining more biological samples, we use nonparametric bootstrap, a random subsampling method with replacement, to mimic the process of generating more data under the same conditions, i.e., the same level of cell engraftment. Bootstrap provides an alternative to the otherwise infeasible task of obtaining more data for the experiment: We first divide the collections of points on the SDPH diagrams into two balanced sets as training and test sets. Then, across 50 bootstrap samples on the training set, we derive the mean of the optimized parameters. For model evaluation, we evaluate the derived model performance across 50 bootstrap samples on the test set and the results are summarized in the following section.

### v2

@@ 4.2 Spatial Texture Decomposition via Local Texture Analysis
We first perform a local texture analysis using the measures and critical sizes as defined in Section 4.1, we seek to analyze the localized textures by computing ālocalā SDPH diagrams at equidistant locations throughout the 3D image sample for each bone and determining how the distribution of textures on the bone differ with respect to the progression of the disease.
We focus our analysis on the textural composition of the knee region in femoral bones by extracting local features within 3D balls centered on a regular grid. Each ball has a radius of 40 voxels in the horizontal plane and 1/5 of the sample thickness in the vertical plane, and ball centres are spaced 5 voxels apart. This yields over 22 million overlapping regions across the full dataset of 27 bone samples. The choice of a 40-voxel radius (approximately 80m, since each voxel measures 2m) balances biological relevance and computational feasibility. At this scale, each ball typically captures multiple sinusoidal capillaries (10ā20m in diameter), along with surrounding tissue and cavities, allowing for informative characterization of local morphology. Smaller radii risk producing sparse or uninformative topological features, while larger radii may over-smooth and obscure important heterogeneity.
To compute SDPH diagrams, we first divide each bone image into 8 coarse subregions over the knee, which support distributed processing and spatial stratification. Within each subregion, we extract SDPH diagrams for each individual ball by including only those topological features whose birth and death points lie entirely within the ball. Each ball-level diagram provides a localized summary of vessel structure. We remove points with persistence less than 0.5 (roughly 0.25m), as these are likely to correspond to imaging noise, given the scale of the vessels of interest (10ā20m).
@@
After normalizing the localized feature vectors, we further group these feature vectors into clusters comprising similar textures. Our dominant approach utilizes the -means algorithm (MacQueen etĀ al.,, 1967) in clustering the locations into groups. The -means algorithm is an iterative method to partition the dataset into clusters of similar features. By initializing on some random centroids, then iteratively assigning data points to their nearest centroid, then updating the centroids with the means of the assigned data points, we converge to stable centroids and cluster assignments for the data. We also considered other approaches using alternative clustering techniques, such as Gaussian mixture model clustering (Reynolds,, 2009) and clustering for large applications (CLARA, Kaufman & Rousseeuw, (1990)) which yielded similar results. These results can be found in Section B of the Supplementary Material.
@@
Figure 7 shows the distributions of groups of different textures in the knee regions of the bone for the different samples when the number of clusters is set to , represented in the figures by the different colors projected back onto their locations in the bone image. The projected distribution for the long regions of the bone is available in the Supplementary Material. From Figure 7, we observe that the CTRL samples consist of largely the white texture, which we arbitrarily assign to be texture A, whereas other samples with higher levels of engraftment contain more of the other two dark and light blue colored-textures, which we denoted as textures B and C. We also identify a transition in the ratio of different textures across the samples. Figure 8 shows the varying percentage decomposition into A, B, C textures for the knee region of each bone sample.
@@
Figure 9 shows the percentage decomposition of the 3 textures projected onto two dimensions for visualization using principal component analysis (PCA), a dimensionality reduction technique which seeks to emphasize the significant features of the data while retaining as much variation of the data as possible. Based on this, we can approximately divide the samples into 3 groups, which we can see highlighted in the figure in blue, orange, and yellow. It is important to note that this grouping is not unique to this specific combination of and -means clustering algorithm; we find that across a range of values of and both the GMM and CLARA clustering algorithms, we repeatedly observe clear divisions between the outlined Phase I and II samples. On the other hand, Phase O samples may appear distinct to the other two groups or merged with the other groups for some combinations of and algorithms, however the occurrence of the latter is the rarer of the two. Certain samples may violate the grouping, for example, for certain combinations, P1 sample at 40% may appear closer to Phase I samples than Phase II samples. Further supporting evidence is shown in Section B of the Supplementary Material.
@@
Overall, based on the percentage texture decomposition, we allocate the knee samples into 3 morphological stagesāPhases O, I, and IIāas follows:
ā¢
@@
Returning to Figure 8, we find that, as previously noted, the dominant texture for Phase O samples is texture A, constituting on average over 60% of the texture decompositionāa considerably higher proportion than that found within the other phases. In contrast, for Phase I samples, the dominance shifts from texture A to texture B, with the samples containing on average between 45% to 55% of texture B. Lastly, for Phase II samples, we observe on average a more balanced distribution between the textures, with a similar proportion for each texture. Morphologically, texture A is characterized by thicker vessels compared to textures B and C, whereas texture B features lower values than the other textures and smaller interstitial gaps between the vessels. In contrast, texture C features much larger interstitial gaps between the vessels than found in the other two textures. Further analysis of individual textures can be found in Song, (2023).
Furthermore, we find a correspondence between the phase allocations and the levels of engraftment. By assignment, we imposed that Phase O be solely associated with the CTRL samples (in the ācontrolā stage with no engraftment), while Phase I is associated with samples with low percentages of engraftment, mostly varying between 1% and 25% of engraftment; we describe this as the early stages of engraftment. In comparison, Phase II contains samples with higher percentages of engraftment, varying between 40% and 90%; we associate this with late stages of engraftment. The exceptions to this trend are the two samples MNC at 53% and P1 at 67%, whose estimated levels of engraftment may differ to the actual levels of engraftment due to the estimation process described previously.
@@ 4.3 Spatial Texture via Global Texture Analysis
In parallel, we performed a global analysis of the textures by computing SDPH on the entire knee regions of the bone samples and observed commonalities in clustering behavior with the phase groupings from the local texture analysis. To visualize the large collection of birthādeath points in persistence diagrams computed for each sample, we compute an approximate density in the form of a Gaussian kernel density estimate. In particular, on each point in the diagram we applied the Gaussian kernel, with , weighted by their persistence, that is, the difference between the birth and death times, similar to the approaches in Maroulas etĀ al., (2019) and Moon etĀ al., (2020). Figure 10 shows an example of such Gaussian kernel density estimates (KDEs) computed for the persistence diagrams of the first CTRL sample plotted as contoured heatmaps.
@@
We observe several different types of transitions in density in the different quadrants of the persistence diagrams as the samples vary from different types and levels of engraftments. In particular, Table 1 shows a selection of persistence-weighted heatmaps of the density estimates for the PH1 NW quadrants of the SDPH diagrams of the knee samples. We observe increasing variation in values for higher levels of engraftment and from Phase I samples to Phase II samples. This variation suggests that the vascular loops dilate towards later stages of engraftment. Also, moving from Phase O to Phase II samples, we observe a shift in the range of values towards lower values of , suggesting thinner vessels and perhaps thinner vessel walls and greater permeability related to later stages of AML.
@@
Another interesting feature is the spread of the points in the PH2 NE quadrant of the persistence diagrams. We note from Table 2 that in general, the samples with higher levels of engraftment show greater variation in the distribution of the points in the direction parallel to the diagonal. Other changes can be similarly found in the other quadrants transitioning through the engraftment levels and phases. The full tables of density heatmaps for the knee and long regions are available within Section C of the Supplementary Material.
-2cm
Phase O
Phase I
Phase II
-2cm
Phase O
Phase I
Phase II
To explore whether the morphological distinctions identified through local texture analysis are also reflected at a global level, we perform hierarchical clustering on the Gaussian KDEs derived from the persistence diagrams. We apply standard agglomerative hierarchical clustering using 4 different linkage strategiesācomplete, average, single, and Wardāand compute a consensus over both linkage methods and bootstrap resamples. This helps mitigate the sensitivity of hierarchical clustering to specific algorithmic choices and yields a more robust global structure across samples. The resulting consensus dendrograms provide global, density-based visualizations of the morphological similarity across samples. The optimal number of clusters, set to 4, was determined via the elbow method based on the within-cluster sum of squares at each number of clusters.
The hierarchical tree computed from PH1 KDEs with bootstrap consensus is shown in FigureĀ 11, and the tree computed using consensus over linkage methods is shown in FigureĀ 12. Given the optimal cluster count of 4, we cut the trees at approximate heights of 0.8 and 1.75, respectively, resulting in 4 major branches. These clusters largely align with the local phase classifications: 3 clusters consisting uniquely of solely Phase O, I and II samples, respectively, and a fourth cluster consisting of mostly Phase I samples, with two exceptionsāCTRL sample at 0% (Phase O) and P1 sample at 40% (Phase II)āsuggesting that these samples share intermediate global texture patterns. Similar consensus hierarchical tree for PH2 can be found in Section D of the Supplementary Materials and supports a similar finding.
To quantify the alignment between the global clustering and local phase groupings, we compute several external clustering evaluation metrics: Adjusted Rand Index (ARI), Normalized Mutual Information (NMI), Normalized Variation of Information (VI), and normalized split/join distance. These results, summarized in TableĀ 3, indicate moderate to strong agreement, particularly for the Phase I and Phase II samples.
Metric | PH Region | Ward | Average | Complete | Single |
ARI | PH0 | -0.015 | 0.023 | -0.009 | 0.071 |
PH1 | 0.626 | 0.626 | 0.502 | 0.673 | |
PH2 | 0.463 | 0.529 | 0.502 | 0.498 | |
NMI | PH0 | 0.131 | 0.157 | 0.137 | 0.325 |
PH1 | 0.761 | 0.761 | 0.638 | 0.716 | |
PH2 | 0.552 | 0.614 | 0.638 | 0.628 | |
VI | PH0 | 0.900 | 0.812 | 0.902 | 0.729 |
PH1 | 0.244 | 0.244 | 0.374 | 0.303 | |
PH2 | 0.464 | 0.399 | 0.374 | 0.431 | |
Split/Join | PH0 | 0.083 | 0.074 | 0.085 | 0.066 |
PH1 | 0.020 | 0.020 | 0.031 | 0.017 | |
PH2 | 0.040 | 0.031 | 0.031 | 0.034 |
@@ 5 Modeling Textural Changes with Gaussian Mixture Models on SDPH Diagrams
In Section 4, we have shown that associations can be made between the SDPH diagrams and morphological features in the bone samples. This indicates that SDPH diagrams can quantify structural changes in bone marrow vasculature in an interpretable manner that is consistent with biological reasoning. Using localized analysis of the images, we assigned the samples into one of 3 morphologically-derived phasesāO, I, and IIāwhich approximately correspond to the control, early, and late stages of engraftment. Further, by computing the global SDPH over the knee regions and approximating the point distributions using a Gaussian kernel density estimate, we confirmed concordance with the phase allocations. We interpreted the transitions in the density patterns within the SDPH diagrams across phases as reflecting biological progression consistent with increasing levels of engraftment. To capture this behavior, we turn to our second contribution of this paper and propose a family of phase-dependent Gaussian mixture models (GMMs) to represent the common distributions observed for samples in each phase. This approach aligns with prior work on the statistical modeling of persistence diagrams, notably the Bayesian framework of Maroulas etĀ al., (2020), but differs in its focus on interpretable, low-dimensional mixtures tailored to biologically relevant, high-persistence features.
@@ 5.1 Gaussian Mixture Models
Gaussian mixture models have been widely used for unsupervised clustering and density estimation (Reynolds,, 2009); the models assume the data to be generated from a weighted mixture of Gaussian distributions which are parameterized by some unknown means and variances. Where the -means algorithm is considered a hard clustering technique, GMMs allow for soft clustering, where cluster assignments are associated with probabilities, allowing the model to capture more complex patterns in the data distribution. It is also common for the parameters of the GMM to depend on external parameters, such as time for applications in finance (Zhang etĀ al.,, 2020) and engineering (Masāud etĀ al.,, 2021), allowing for context to be integrated into the model.
Due to the discrete nature of our samples and uncertainty in terms of the levels of engraftment, we introduce instead a family of phase-dependent GMMs to approximate our samplesā SDPH diagrams. For each point in the SDPH diagram for each sample, we relabel the indices, so that points from samples of the same phase are viewed as a collective: Let for denote the collection of birthādeath points from the SDPH diagrams of samples within phase , where denotes the number of points in this phase and references the individual points in this collection. In addition, each point is weighted by its persistence computed as . Let be the size of the mixture for each phase, then the models are given by
@@
ExpectationāMaximization (EM) Algorithm. To estimate the parameters , we use an iterative expectationāmaximization (EM) algorithm, which can be broken down into two steps: An E-step, computing the probability of belonging to each Gaussian distribution,
@@
We determine the size of the mixture model through trade-offs between the model performance evaluated using the Bayesian information criterion (BIC), the computational time, and the generalizability of the model across SDPH diagrams of all samples in the same phase. BIC is chosen for the application since it poses heavier penalties on the model complexity compared to other criteria, such as the Akaike information criterion (AIC), and is ideal for our application.
@@
Since we have no further way of generating more images from each engrafted sample or obtaining more biological samples, we use nonparametric bootstrap, a random subsampling method with replacement, to mimic the process of generating more data under the same conditions, i.e., the same level of cell engraftment.
A Continuous Regression-Based Model. While our approach models disease progression using 3 discrete phases, a continuous regression-based framework could, in principle, offer a more detailed view of structural change. A natural candidate is a ātimeā-varying Gaussian Mixture Model, where the component weights , means , and covariances vary with engraftment level :
However, this approach faces both biological and modeling limitations. Biologically, engraftment levels are measured at whole-body scale and may not reliably reflect local bone marrow states. In addition, inter-animal variability and different AML subtypes introduces further heterogeneity that complicates modeling. From a statistical perspective, the sparsity and repetition of time values, along with label noise and numerical instability, limit predictive performance. These challenges make the regression model suboptimal for our dataset under study. We retain the phase-dependent formulation for its robustness and interpretability, and refer to Section E of the Supplementary Material for additional details and results for the proposed continuous model.
5.2 Model Interpretation
In this section, we fit our models to the full dataset to interpret the morphological structure characteristics of each phase. By examining the fitted distributions within each quadrant individually, we are able to identify the dominant structural features and highlight how these evolve across phases. This enables us to characterize the progression of morphological changes from Phase O through to Phase II, offering a global summary of the vascular remodeling associated with AML progression.
PH1 NW. We begin our evaluation by studying the NW quadrant which characterizes the number of loops, their thickness, and sizes, and was seen to give rise to distinguishing features in the samplesā estimated density between the different phases. On NW samples, taking into account the trade-off between the model fit, computational time, and ensuring that the model is generic across all the samples within one phase, we found that the optimal number of components for the Phase O model to be 3, then increasing the number of components for each consecutive phase by 1 was found to distinguish the models sufficiently as well as allow good fits to the samples within each phase. Figure 13 shows the GMMs for each phase and interpreting from the parameter of the GMMs, we found that compared to the CTRL samples, the Phase I samples tend to contain thinner and more compact vessels in the early stages of AML engraftment. Transitioning from Phase I to II, towards later stages of engraftment, we observe instead a shift in density towards the northwest direction in the diagram, in favor of higher values of and sizes, indicative of thicker and wider loops formed by vessels in the cavity, although not as much as those found in Phase O.
PH1 NE. We next study the neighboring quadrant PH1 NE. From the transitions between the Phase O, I, and II models for NE shown in Figure 14, we find that compared to the unaffected bone marrow vasculature in the CTRL samples, samples in the early stages of AML engraftment have more compact vasculature. We observe a higher probability around lower values of both and critical sizes, suggesting tighter spacing between vessels in general, both in the narrowing between loops and loops formed by vessels. Then, transitioning from the Phase I model to the Phase II model, we observe the region of high probability moving towards higher and values, suggesting a dilation of the underlying vessel structure.
PH2 NE. Evaluating our model on the NE quadrant. Between the Phase O and I models, we again find a shift in the region of high probability towards lower values of and , indicating smaller loops and cavities formed by vessels. This suggests a more compact vasculature at the onset of angiogenesis. As we transition from the Phase I to the Phase II model, we find that the distribution shifts towards larger values of and , indicating that the vasculature becomes more sparse. Figure 15 shows the composition of the Gaussian mixture model and densities plotted as heatmaps.
The collection of points in the NW quadrant can be disregarded since they arise as a complication in the piecewise computation of SDPH along the bone images. In theory, there should only be one point with the quadrant for each sample representing the entire connected vessel structure. The points in the NW can also be disregarded since it contains only a few points which should never, or only very rarely, occur since they correspond to bubbles of extravascular space trapped inside the lumen and may only appear due to noise. The evaluation for the SW and SW quadrant are available in Section F of the Supplementary Material as the type of characteristics described within these quadrants are not distinct between the phases.